# Cartesian Nodal Lines and Magnetic Kramers Weyl Nodes in Spin-Split Antiferromagnets

### v1

@@ Abstract
Introduction.—Momentum-dependent spin splitting (MDSS) serves as a pivotal driving force behind the emergence of non-trivial quantum geometry and the realization of a diverse range of topological phases. Spin-orbit coupling (SOC) in noncentrosymmetric systems is a fundamental mechanism responsible for this phenomenon and has sparked extensive research over the past two decades [1, 2, 3, 4, 5, 6]. Due to its time-reversal ($\mathcal{T}$)-even and inversion ($\mathcal{P}$)-odd nature, SOC-induced MDSS exhibits intrinsic nodes at time-reversal-invariant momenta. These nodes manifest as band degeneracies within the band structure and underpin various topological phases [7, 8, 9, 10, 11, 12].
@@
Exchange interaction is another fundamental mechanism for spin splitting. Recent discoveries have revealed that even in antiferromagnets with zero-net magnetization, spin splitting can exhibit significant strength and momentum dependence [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], provided that the system lacks $\mathcal{PT}$ symmetry or $\mathcal{T}\tau$ symmetry, where $\tau$ represents a translation operation. Notably, the exchange-interaction-induced MDSS in spin-split antiferromagnets also features nodes, resulting in symmetry-enforced band degeneracies. For example, this is seen in spin-split antiferromagnets with collinear magnetic moments, also known as altermagnets [21, 27]. These materials have garnered significant attention due to their unique spin-split band structures [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] and a wide range of intriguing phenomena they host [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. As spin is conserved for a collinear magnetic order, the nodes in the MDSS form nodal surfaces in three dimensions (3D) and nodal lines in two dimensions (2D) in the absence of SOC. However, these band degeneracies are characterized by a codimension of $d_{c}=1$ and do not lead to topological boundary states [75].
@@
In search of band degeneracies with nontrivial properties, we are led to examine spin-split antiferromagnets with noncollinear magnetic moments [14, 16, 23]. For these materials, the spin conservation is intrinsically broken and the nodes in the exchange-interaction-induced MDSS are severely constrained by crystal symmetries. By analyzing additional constraints imposed by these symmetries and the interplay of SOC and exchange interaction, we uncover two new classes of band degeneracies with fascinating properties, which we refer to as the Cartesian nodal lines (CNLs) and magnetic Kramers Weyl nodes (MKWNs) respectively.
@@
Without SOC, band degeneracies in 3D spin-split noncollinear antiferromagnets generally take the form of nodal lines with a codimension $d_{c}=2$ [76]. Since the $\mathcal{PT}$ symmetry is absent, these nodal lines are protected by mirror symmetry and are confined to mirror planes. Additional crystal symmetries further constrain them to intersect and form a structure resembling the Cartesian coordinate system, hence the name Cartesian nodal lines. Distinguished from other nodal-line structures protected by $\mathcal{PT}$ symmetry or chiral symmetry [77, 78, 79, 80, 81, 82, 83, 84, 85], these CNLs give rise to not only unique Berry curvature distributions but also topological surface states with unconventional spin textures. In the presence of SOC, these CNLs undergo a transition into Weyl nodes with $d_{c}=3$. Notably, some of these Weyl nodes are pinned at specific time-reversal invariant momenta. Reminiscent of the Kramers Weyl nodes protected by time-reversal symmetry in chiral crystals [99, 100], these Weyl nodes are thus dubbed magnetic Kramers Weyl nodes. The existence of these band degeneracies has far-reaching consequences, as they can lead to, under suitable conditions, strong or even quantized anomalous Hall effects as well as quantized circular photogalvanic effects.
@@
Cartesian nodal lines.—Constrained by symmetry, the exchange-interaction-induced MDSS can be viewed as an order parameter analogous to the superconducting pairing [86], and can be classified by the irreducible representations of symmetry groups [76]. While the discussed physics in this paper is general, we focus on a cubic-lattice antiferromagnet within the $D_{4h}$ point group and a specific MDSS described by the $B_{2g}^{-}$ irreducible representation for illustration (see Table I in Ref.[76]). Accordingly, the minimal effective tight-binding Hamiltonian describing the spin-split band structure is given by $\hat{H}=\sum_{k}c_{\boldsymbol{k}}^{\dagger}\mathcal{H}(\boldsymbol{k})c_{\boldsymbol{k}}$, where $c_{\boldsymbol{k}}^{\dagger}=(c_{\boldsymbol{k}\uparrow}^{\dagger},c_{\boldsymbol{k}\downarrow}^{\dagger})$ and
@@
Here $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ denotes the vector of Pauli matrices, and $\sigma_{0}$ represents the identity matrix. The first term with $\varepsilon_{0}(\boldsymbol{k})=-t(\cos k_{x}+\cos k_{y})-t_{z}\cos k_{z}$ refers to the kinetic energy, the second term with $\boldsymbol{l}(\boldsymbol{k})=(\sin k_{x},\sin k_{y},\sin k_{z})$ is the SOC, and the last term with $\boldsymbol{m}(\boldsymbol{k})=[-\sin k_{x}\sin k_{z},\sin k_{y}\sin k_{z},\eta(\cos k_{x}-\cos k_{y})]$ accounts for the magnetic exchange field [76]. The parameter $\eta$ is introduced to characterize the anisotropy and is typically nonzero. For notational simplicity, we set the lattice constants to unity throughout this work.
Let us first focus on the MDSS induced solely by the magnetic exchange field, i.e., $\lambda_{\rm so}=0$. In this case, the Hamiltonian has inversion symmetry, three mirror symmetries ($\mathcal{M}_{z}$, $\mathcal{M}_{xy}$, $\mathcal{M}_{\bar{x}y}$), and rotation/time-reversal combinational symmetry ($C_{4z}\mathcal{T}$, $C_{2x}\mathcal{T}$, $C_{2y}\mathcal{T}$), where $C_{na}$ represents a $2\pi/n$ rotation about the $a$ axis. These $C_{na}\mathcal{T}$ symmetries ensure a zero net magnetization. The corresponding energy spectra are
$E_{\pm}(\boldsymbol{k})$ $=$ $\varepsilon_{0}(\boldsymbol{k})\pm\lambda_{\rm M}\left[\sin^{2}k_{z}(\sin^{2}k_{x}+\sin^{2}k_{y})\right.$ (2)
$\left.+\eta^{2}(\cos k_{x}-\cos k_{y})^{2}\right]^{1/2}.$
@@
It is clear from the above spectra that band degeneracies appear at these positions: (1) along the momentum lines satisfying $|k_{x}|=|k_{y}|$ within the $k_{z}=0/\pi$ planes, and (2) along the high-symmetry $k_{z}$ lines passing through $(k_{x},k_{y})=(0/\pi,0/\pi)$, as illustrated in Fig. 1(a). These nodal lines are protected by the three mirror symmetries, and intersect orthogonally at the four $C_{4z}\mathcal{T}$-invariant momenta within the Brillouin zone, i.e., $(0,0,0/\pi)$ and $(\pi,\pi,0/\pi)$. At each intersection, the nodal-line structure is analogous to the Cartesian coordinate system, suggesting to us the name CNL. The CNLs can be considered a distinct class of crossed $Z_{3}$ nodal nets; however, their origin and properties differ significantly from those found in nonmagnetic [87] and altermagnetic materials [88].
Figure 1: (a) Solid and dashed orange lines represent nodal lines, which form a structure analogous to the Cartesian coordinate system at each intersection. (b) Distribution of the Berry curvature in the momentum plane with $k_{z}=0.1$. (c) Energy spectra along a path in the surface Brillouin zone, with open boundary conditions in the $[\bar{1}10]$ direction. High-symmetry points in the surface Brillouin zone are shown in the inset. (d) Black arrows and gradient colors jointly depict the spin polarizations of the surface states on the right $(\bar{1}10)$ surface. The green rings indicate the projection of the bulk Fermi surface, with the chemical potential fixed at $\mu=0.1$. Common parameters are $t=t_{z}=\lambda_{\rm so}=0$, and $\lambda_{\rm M}=\eta=1$.
@@
In nonmagnetic materials with negligible SOC, nodal lines are typically protected by $\mathcal{PT}$ symmetry [89], which ensures the vanishing of the Berry curvature [90]. Here, however, the $\mathcal{PT}$ symmetry is broken, and the CNLs are instead protected by mirror symmetry. Consequently, finite Berry curvature is not only permitted but becomes divergently large near the band degeneracy, as illustrated in Fig.1(b). The distribution of the Berry curvature respects the symmetries of the point group, resulting in an exact cancellation of the Hall conductivity when the anomalous Hall effect is considered [91]. Nevertheless, its divergent nature near the band degeneracy implies that an external perturbation, which disrupts the symmetry-enforced cancellation, can induce a strong anomalous Hall effect. This will be demonstrated in detail later.
@@
A characteristic of nodal lines is the emergence of dispersionless topological surface states when chiral symmetry is present [92]. These topological surface states exhibit fixed spin polarizations, as they simultaneously serve as eigenstates of the chiral symmetry operator—a constant unitary and Hermitian operator that anticommutes with the Hamiltonian [93]. Intriguingly, we discover that in this system, topological surface flat bands exist even in the absence of chiral symmetry, provided that the term $\varepsilon_{0}(\boldsymbol{k})\sigma_{0}$, which has no impact on topology, is omitted from the Hamiltonian. In Fig.1(c), the energy spectrum for a system with open boundary conditions along the $[\bar{1}10]$ direction is plotted along a specific high-symmetry path in the surface Brillouin zone. The existence of surface flat bands at zero energy is clearly visable. Furthermore, the region supporting topological surface states spans the whole surface Brillouin zone, except for these four high-symmetry lines where the bulk spectrum is gapless. This is because the CNLs form a network and their projection along the $[\bar{1}10]$ direction overlaps with the four high-symmetry lines of the surface Brillouin zone. By analyzing the spin textures of these topological surface states, we observe that their spin polarizations are not fixed but instead exhibit strong momentum dependence, as illustrated in Fig.1(d). Interestingly, the spin textures in the four quadrants of the surface Brillouin are symmetry-related and form a plaid-like configuration. This distinctive pattern of spin textures can serve as a definitive signature for identifying the presence of CNLs in experiments.
How do we understand the existence of topological surface flat bands in the absence of chiral symmetry, as well as the distinctive pattern of the spin textures? We find that these counterintuitive results can be attributed to the existence of subchiral symmetry in this system. Subchiral symmetry, a concept introduced in Ref.[94], generalizes the notion of chiral symmetry and has proven to be widely applicable in understanding the properties of topological boundary states [95, 96]. To see the connection between the surface flat bands and the subchiral symmetry, we first choose an appropriate coordinate system to calculate the surface states. For these surface states on the $(\bar{1}10)$ surfaces, the most convenient is one that rotates $\pi/4$ around the $z$-axis relative to the original coordinate. In the new coordinate system, the momentum ${\boldsymbol{k}}=(k_{1},k_{2},k_{z})$ and the Pauli-matrix vector $\boldsymbol{\sigma}=(\sigma_{-},\sigma_{+},\sigma_{z})$ , where $k_{1}=\frac{k_{x}-k_{y}}{\sqrt{2}}$, $k_{2}=\frac{k_{x}+k_{y}}{\sqrt{2}}$ [see Fig.1(a)], $\sigma_{-}=\frac{\sigma_{x}-\sigma_{y}}{\sqrt{2}}$ and $\sigma_{+}=\frac{\sigma_{x}+\sigma_{y}}{\sqrt{2}}$. The Hamiltonian without SOC, denoted by $\mathcal{H}_{\rm CNL}(\boldsymbol{k})$, becomes (see details in the supplemental material (SM) [97])
$\mathcal{H}_{\rm CNL}(\boldsymbol{k})$ $=-\sqrt{2}\lambda_{\rm M}\sin k_{z}\cos\frac{k_{1}}{\sqrt{2}}\sin\frac{k_{2}}{\sqrt{2}}\ \sigma_{-}$
-- $\lambda_{\rm M}\Lambda(k_{2},k_{z})\sin\frac{k_{1}}{\sqrt{2}}\left(\sin\theta\ \sigma_{z}+\cos\theta\ \sigma_{+}\right),$ (3)
where $\theta=\theta(k_{2},k_{z})=\arg(\sqrt{2}\sin k_{z}\cos\frac{k_{2}}{\sqrt{2}}+i2\eta\sin\frac{k_{2}}{\sqrt{2}})$ and $\Lambda(k_{2},k_{z})=\sqrt{2\sin^{2}k_{z}\cos^{2}\frac{k_{2}}{2}+4\eta^{2}\sin^{2}\frac{k_{2}}{\sqrt{2}}}$. We have dropped the term $\varepsilon_{0}(\boldsymbol{k})\sigma_{0}$ since it has no impact on the band topology. Although $\mathcal{H}_{\rm CNL}(\boldsymbol{k})$ lacks chiral symmetry, it has subchiral symmetry since it anticommutes with the following momentum-dependent unitary and Hermitian operator,
$\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}.$ (4)
For surface states on the $(\bar{1}10)$ surfaces, $k_{2}$ and $k_{z}$ are good quantum numbers and hence can be viewed as parameters. For fixed $k_{2}$ and $k_{z}$, $\mathcal{S}(k_{2},k_{z})$ plays the exact role as a chiral symmetry operator. Therefore, the zero-energy topological surface states presented in Fig.1(c) can be characterized by the following winding number [98],
[…]
The spectrum along the $k_{1}$ direction exhibits insulating behavior as long as $k_{2}\neq\{0,\sqrt{2}\pi\}$ and $k_{z}\neq\{0,\pi\}$. In the region where a gap exists, we observe that $W(k_{2},k_{z})=\text{sgn}(k_{2}k_{z})$, where $\text{sgn}(\cdot)$ denotes the sign function. The nonzero winding number indicates the presence of zero-energy states at the boundary, which collectively form the surface flat bands.
Since $\{\mathcal{S}(k_{2},k_{z}),\mathcal{H}_{\rm CNL}(\boldsymbol{k})\}=0$, the zero-energy surface states simultaneously serve as eigenstates of the subchiral symmetry operator $\mathcal{S}(k_{2},k_{z})$. Consequently, their spin textures are straightforward to determine. The results are [97]
@@
where $\beta=1$ $(-1)$ refers to the left (right) surface, indicating that the spin textures on opposing surfaces are of opposite orientation. These analytical results provide an intuitive and consistent explanation for the numerical findings presented in Figs.1(c) and 1(d).

### v3

@@ Abstract
Introduction.—Momentum-dependent spin splitting (MDSS) serves as a pivotal driving force behind the emergence of non-trivial quantum geometry and the realization of a diverse range of topological phases. Spin-orbit coupling (SOC) in noncentrosymmetric systems is a fundamental mechanism responsible for this phenomenon and has sparked extensive research over the past two decades [1, 2, 3, 4, 5, 6]. Due to its time-reversal ($\mathcal{T}$)-even and inversion ($\mathcal{P}$)-odd nature, SOC-induced MDSS exhibits intrinsic nodes at time-reversal-invariant momentums (TRIMs). These nodes manifest as band degeneracies within the band structure and underpin various topological phases [7, 8, 9, 10, 11, 12].
@@
Exchange interaction is another fundamental mechanism for spin splitting. Recent discoveries have revealed that even in antiferromagnets with zero-net magnetization, spin splitting can exhibit significant strength and momentum dependence [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], provided that the system lacks $\mathcal{PT}$ symmetry or $\mathcal{T}\tau$ symmetry, where $\tau$ represents a translation operation. Notably, the exchange-interaction-induced MDSS in spin-split antiferromagnets also features nodes, resulting in symmetry-enforced band degeneracies. For example, this is seen in spin-split antiferromagnets with collinear magnetic moments, also known as altermagnets [21, 27]. These materials have garnered significant attention due to their unique spin-split band structures [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] and a wide range of intriguing phenomena they host [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. As spin is conserved for a collinear magnetic order, the nodes in the MDSS form nodal surfaces in three dimensions (3D) and nodal lines in two dimensions (2D) in the absence of SOC. However, these band degeneracies are characterized by a codimension of $d_{c}=1$ and generally do not lead to topological boundary states [76].
@@
In search of band degeneracies with nontrivial properties, we are led to examine spin-split antiferromagnets with noncollinear magnetic moments [14, 16, 23]. For these materials, the spin conservation is intrinsically broken and the nodes in the exchange-interaction-induced MDSS are severely constrained by crystal symmetries. By analyzing additional constraints imposed by these symmetries and the interplay of SOC and exchange interaction, we uncover two unique forms of band degeneracies with fascinating properties, which we refer to as the Cartesian nodal lines (CNLs) and magnetic Kramers Weyl nodes (MKWNs) respectively.
@@
Without SOC, band degeneracies in 3D spin-split noncollinear antiferromagnets generally take the form of nodal lines with a codimension $d_{c}=2$ [77]. Since the $\mathcal{PT}$ symmetry is absent, these nodal lines are protected by mirror symmetry and are confined to mirror planes. Additional crystal symmetries further constrain them to intersect and form a structure resembling the Cartesian coordinate system, hence the name Cartesian nodal lines. Distinguished from other nodal-line structures protected by $\mathcal{PT}$ symmetry or chiral symmetry [78, 79, 80, 81, 82, 83, 84, 85, 86, 87], these CNLs give rise to not only unique Berry curvature distributions but also topological surface states with unconventional spin textures. In the presence of SOC, these CNLs undergo a transition into Weyl nodes with $d_{c}=3$. Notably, some of these Weyl nodes are pinned at specific TRIMs. Reminiscent of the Kramers Weyl nodes protected by time-reversal symmetry in chiral crystals [88, 89], these Weyl nodes are thus dubbed magnetic Kramers Weyl nodes. The existence of these band degeneracies has far-reaching consequences, as they can lead to, under suitable conditions, strong or even quantized anomalous Hall effects as well as quantized circular photogalvanic effects.
Figure 1: Schematic of the lattice structure. White (blue) spheres represent nonmagnetic (magnetic) atoms, while red arrows indicate the local magnetic moments. [(a)] presents the magnetic background surrounding a representative nonmagnetic atom; the configurations around other nonmagnetic atoms are identical and thus omitted for clarity. (b) Schematic illustration of the Wigner-Seitz unit cell. The spin (red) and real-space (black) coordinate conventions are the same as in [(a)]. (c-e) Schematics of the spin-dependent hoppings of the form $\boldsymbol{S}\cdot\boldsymbol{\sigma}$ in the $xy$, xzxz, and $yz$ planes, respectively. Dashed black arrows indicate hopping directions, and the corresponding amplitudes are shown in light-blue rounded rectangles. Panels [(d,e)] share the same spin-coordinate convention as [(c)].
@@
Cartesian nodal lines.—Constrained by symmetry, the exchange-interaction-induced MDSS can be viewed as an order parameter analogous to the superconducting pairing [90], and can be classified by the irreducible representations of symmetry groups [77]. While the physics discussed here is general, we focus on a tetragonal-lattice antiferromagnet belonging to the $D_{2h}$ magnetic point group of the $D_{4h}$ point group (magnetic group $D_{4h}$ ($D_{2h}$) in the $G(H)$ convention). A schematic of the lattice structure is shown in Fig. 1. We assume that the electrons associated with magnetic atoms are strongly localized, and the itinerant electrons, which are responsible for the system’s transport behaviors, originate from nonmagnetic atoms. The magnetic atoms act as a magnetic background and influence the itinerant electrons as spin-dependent potentials. When the itinerant electrons hop along paths intersecting these magnetic atoms, their hoppings become spin-dependent as an electron with spin parallel or antiparallel to the local magnetic moment experiences a different tunneling probability. As a result, the tight-binding Hamiltonian describing the spin-split band structure of itinerant electrons takes the form $\hat{H}=\sum_{k}c_{\boldsymbol{k}}^{\dagger}\mathcal{H}(\boldsymbol{k})c_{\boldsymbol{k}}$, where $c_{\boldsymbol{k}}^{\dagger}=(c_{\boldsymbol{k}\uparrow}^{\dagger},c_{\boldsymbol{k}\downarrow}^{\dagger})$ and
@@
Here, $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ denotes the vector of Pauli matrices, and $\sigma_{0}$ represents the identity matrix. The first term, with $\varepsilon_{0}(\boldsymbol{k})=-t(\cos k_{x}+\cos k_{y})-t_{z}\cos k_{z}$ refers to the kinetic energy. The second term represents spin-orbit coupling (SOC), $\boldsymbol{l}(\boldsymbol{k})=(\sin k_{x},\sin k_{y},\sin k_{z})$, which is forbidden by the spatial symmetries (including inversion and mirror symmetries) of the lattice in Fig. 1, but can emerge once these symmetries are broken, for instance by lattice distortion. The last term accounts for the magnetic exchange field [77] with $\boldsymbol{m}(\boldsymbol{k})=[-\sin k_{x}\sin k_{z},\sin k_{y}\sin k_{z},\eta(\cos k_{x}-\cos k_{y})]$, where the parameter $\eta$ characterizes the magnetic anisotropy and is typically nonzero. For notational simplicity, we set the lattice constants to unity throughout this work.
@@
We begin by examining the MDSS arising purely from the magnetic exchange field in the lattice shown in Fig. 1, where SOC is forbidden. In this case, the system respects inversion symmetry, three mirror symmetries ($\mathcal{M}_{z}$, $\mathcal{M}_{xy}$, $\mathcal{M}_{\bar{x}y}$), and combined rotation-time-reversal symmetries ($C_{4z}\mathcal{T}$, $C_{2x}\mathcal{T}$, $C_{2y}\mathcal{T}$), where $C_{na}$ represents a $2\pi/n$ rotation about the $a$ axis. These $C_{na}\mathcal{T}$ symmetries ensure a zero net magnetization. The corresponding energy spectra are $E_{\pm}(\boldsymbol{k})=\varepsilon_{0}(\boldsymbol{k})\pm\lambda_{\rm M}\sqrt{\sin^{2}k_{z}(\sin^{2}k_{x}+\sin^{2}k_{y})+\eta^{2}(\cos k_{x}-\cos k_{y})^{2}}$. It is clear from the spectra that band degeneracies appear at these positions: (1) along the momentum lines satisfying $|k_{x}|=|k_{y}|$ within the $k_{z}=0/\pi$ planes, and (2) along the high-symmetry $k_{z}$ lines passing through $(k_{x},k_{y})=(0/\pi,0/\pi)$, as illustrated in Fig. 2(a). Owing to the polar-vector nature of spin—spin components parallel to a mirror plane flip sign, while those perpendicular remain invariant—these nodal lines are protected by the three mirror symmetries. They intersect orthogonally at the four $C_{4z}\mathcal{T}$-invariant momenta within the Brillouin zone, i.e., $(0,0,0/\pi)$ and $(\pi,\pi,0/\pi)$. At each intersection, the nodal-line structure is analogous to the Cartesian coordinate system, motivating the term CNL. While CNLs can be considered a distinct class of crossed $Z_{3}$ nodal nets, their origin and properties differ significantly from those found in nonmagnetic [91] and altermagnetic materials [92].
@@
Unlike nodal-line semimetals protected by $\mathcal{PT}$ symmetry [93], where the Berry curvature is identically zero [94], mirror-protected CNLs permit finite Berry curvature, which becomes divergently large near the band degeneracy, as illustrated in Fig. 2(b). The distribution of the Berry curvature respects the symmetries of the point group, resulting in an exact cancellation of the Hall conductivity when the anomalous Hall effect is considered [95]. Nevertheless, its divergent nature near the band degeneracy implies that an external perturbation, which disrupts the symmetry-enforced cancellation, can induce a strong anomalous Hall effect. This will be demonstrated in detail later.
@@
A characteristic of nodal lines is the emergence of dispersionless topological surface states when chiral symmetry is present [96]. These topological surface states exhibit fixed spin polarizations, as they simultaneously serve as eigenstates of the chiral symmetry operator—a constant unitary and Hermitian operator that anticommutes with the Hamiltonian [97]. Intriguingly, we discover that in this system, topological surface flat bands exist even in the absence of chiral symmetry, provided that the term $\varepsilon_{0}(\boldsymbol{k})\sigma_{0}$, which has no impact on topology, is omitted from the Hamiltonian. In Fig. 2(c), the energy spectrum for a system with open boundary conditions along the $\bar{1}10]$ direction is plotted along a specific high-symmetry path in the surface Brillouin zone. The existence of surface flat bands at zero energy is clearly visible. Furthermore, the region supporting topological surface states spans the whole surface Brillouin zone, except for these four high-symmetry lines where the bulk spectrum is gapless. This is because the CNLs form a network and their projection along the $\bar{1}10]$ direction overlaps with the four high-symmetry lines of the surface Brillouin zone. By analyzing the spin textures of these topological surface states, we find that their spin polarizations are not fixed but instead exhibit strong momentum dependence, as illustrated in Fig. 2(d). Interestingly, the spin textures in the four quadrants of the surface Brillouin zone are symmetry-related and form a plaid-like configuration. This distinctive pattern of spin textures can serve as a definitive signature for identifying the presence of CNLs in experiments.
To understand the existence of topological surface flat bands in the absence of chiral symmetry and the accompanying spin textures, we invoke the concept of subchiral symmetry, as introduced in Ref. [98]. Subchiral symmetry plays the role of chiral symmetry along one-dimensional momentum lines when the perpendicular momentum components are treated as parameters, even though global chiral symmetry is broken. This generalization of chiral symmetry has proven useful in various systems [99, 100]. To make the subchiral symmetry explicit, we rotate the coordinate frame by $\pi/4$ around the $z$-axis. In the rotated frame, the momentum ${\boldsymbol{k}}=(k_{1},k_{2},k_{z})$, where $k_{1}=\frac{k_{x}-k_{y}}{\sqrt{2}}$, $k_{2}=\frac{k_{x}+k_{y}}{\sqrt{2}}$ [see Fig. 2(a)], and the Pauli-matrix vector transforms to $\boldsymbol{\sigma}=(\sigma_{-},\sigma_{+},\sigma_{z})$, where $\sigma_{\pm}=\frac{\sigma_{x}\pm\sigma_{y}}{\sqrt{2}}$. The SOC-free Hamiltonian, denoted by $\mathcal{H}_{\rm CNL}(\boldsymbol{k})$, becomes (see Supplemental information (SI) Sec. I.A for details)
$\mathcal{H}_{\rm CNL}(\boldsymbol{k})$ $=$ $-\sqrt{2}\lambda_{\rm M}\sin k_{z}\cos\frac{k_{1}}{\sqrt{2}}\sin\frac{k_{2}}{\sqrt{2}}\ \sigma_{-}$ (2)
$-\lambda_{\rm M}\Lambda(k_{2},k_{z})\sin\frac{k_{1}}{\sqrt{2}}\sigma_{g}(k_{2},k_{z}),$
where $\sigma_{g}(k_{2},k_{z})=\sin\theta(k_{2},k_{z})\sigma_{z}+\cos\theta(k_{2},k_{z})\sigma_{+}$, $\theta(k_{2},k_{z})=\arg(\sqrt{2}\sin k_{z}\cos\frac{k_{2}}{\sqrt{2}}+i2\eta\sin\frac{k_{2}}{\sqrt{2}})$ and $\Lambda(k_{2},k_{z})=\sqrt{2\sin^{2}k_{z}\cos^{2}\frac{k_{2}}{\sqrt{2}}+4\eta^{2}\sin^{2}\frac{k_{2}}{\sqrt{2}}}$. We have dropped the term $\varepsilon_{0}(\boldsymbol{k})\sigma_{0}$ as it does not affect the band topology. Crucially, $\sigma_{g}(k_{2},k_{z})$ is unitary and anticommutes with $\sigma_{-}$. Hence, there exists a third unitary matrix $\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}$ that anticommutes with both $\sigma_{g}(k_{2},k_{z})$ and $\sigma_{-}$, completing the Pauli-matrices-like Clifford algebra. This matrix $\mathcal{S}(k_{2},k_{z})$ defines the subchiral symmetry operator since $\{\mathcal{S}(k_{2},k_{z}),\mathcal{H}_{\rm CNL}(\boldsymbol{k})\}=0$. For surface states on the $(\bar{1}10)$ surfaces, $k_{2}$ and $k_{z}$ are good quantum numbers and hence can be viewed as parameters. For fixed $k_{2}$ and $k_{z}$, $\mathcal{S}(k_{2},k_{z})$ becomes a constant matrix similar to the operator for chiral symmetry, and accordingly, subchiral symmetry plays the role of chiral symmetry. Therefore, the zero-energy topological surface states presented in Fig. 2(c) can be characterized by the following winding number [101], $W(k_{2},k_{z})=\frac{1}{4\pi i}\int_{-\sqrt{2}\pi}^{\sqrt{2}\pi}dk_{1}{\rm Tr}\left[\mathcal{S}(\mathcal{H}_{\rm CNL})^{-1}\partial_{k_{1}}\mathcal{H}_{\rm CNL}\right]$. The spectrum along the $k_{1}$ direction exhibits insulating behavior as long as $k_{2}\neq\{0,\sqrt{2}\pi\}$ and $k_{z}\neq\{0,\pi\}$. In the gapped region, we find $W(k_{2},k_{z})=\text{sgn}(k_{2}k_{z})$, where $\text{sgn}(\cdot)$ denotes the sign function. The nonzero winding number indicates the presence of zero-energy states at the boundary, which collectively form the surface flat bands. Since these zero-energy surface states simultaneously serve as eigenstates of the subchiral symmetry operator $\mathcal{S}(k_{2},k_{z})$. Their spin textures follow directly: (see Note S1 in SI)
$\braket{\sigma_{z}}(k_{2},k_{z})$ $=\beta{\rm sgn}(k_{2}k_{z})\cos\theta(k_{2},k_{z}),$ (3)
@@
where $\beta=1$ (1)(-1) refers to the left (right) surface, indicating that the spin polarization on opposing surfaces is reversed. These analytical results provide an intuitive and consistent explanation for the numerical observations in Figs. 2(c) and (d).
Figure 2: Cartesian nodal lines and their associated topological properties. (a) Solid and dashed orange lines represent nodal lines, which form a structure analogous to the Cartesian coordinate system at each intersection. (b) Distribution of the Berry curvature in the momentum plane with $k_{z}=0.1$. (c) Energy spectra along a path in the surface Brillouin zone, with open boundary conditions in the $\bar{1}10]$ direction. High-symmetry points in the surface Brillouin zone are shown in the inset. (d) Black arrows and gradient colors jointly depict the spin polarizations of the surface states on the right $(\bar{1}10)$ surface. The yellow network marks the boundary of the zero-energy surface states. Common parameters are $t=t_{z}=\lambda_{\rm so}=0$, and $\lambda_{\rm M}=\eta=1$.
Magnetic Kramers Weyl nodes.—Finite SOC is introduced when inversion and mirror symmetries are broken by symmetry-lowering perturbations, such as strain-induced lattice distortion that gives rise to an asymmetric distribution of the electric potential around the nonmagnetic atoms. In this situation, the magnetic point group is reduced from $D_{4h}$ ($D_{2h}$) to $C_{4}$ ($C_{2}$). Although the system becomes inversion-asymmetric, the essential spin physics can still be accurately captured by the two-band Hamiltonian [88] in Eq. (1). The breaking of these mirror symmetries gaps the previously protected CNLs. However, the surviving $C_{4z}\mathcal{T}$ symmetry mandates the existence of Weyl nodes at the four $C_{4z}\mathcal{T}$-invariant momenta, which coincide with the TRIMs. Since these Weyl nodes are pinned at TRIMs—similar to the Kramers Weyl nodes protected by time-reversal symmetry in chiral crystals[88, 89]—we refer to them as MKWNs to highlight the absence of time-reversal symmetry in this system. Their topological charges $\mathcal{C}_{\boldsymbol{q}}$ are determined by the Chern number $\mathcal{C}=\frac{1}{2\pi}\oint\boldsymbol{\Omega}\cdot d\boldsymbol{S}$, where $\boldsymbol{\Omega}$ represents the Berry curvature, and the integral is performed over a closed surface enclosing the Weyl node located at momentum $\boldsymbol{q}$ [102]. These topological charges are illustrated in the inset located between Figs. 3(a) and (b). We note that magnets can also host a distinct class of Weyl fermions pinned at TRIMs, namely Heesch Weyl fermions [103], which are not protected by an antiunitary symmetry. By contrast, MKWNs require such a symmetry (e.g., $C_{4z}\mathcal{T}$ here) and therefore exhibit a Kramers-like character.